# FlashTex: Fast Relightable Mesh Texturing with LightControlNet

### v1

@@ 5 Experiments
Dataset. As illustrated in Figure 3, we employ Objaverse [10] to render paired data to train our LightControlNet. Objaverse consists of approximately 800k objects, of which we use the names and tags as their text descriptions. We filter out objects with low CLIP similarity [36] to their text descriptions and select around 40k as our training set. Each object is rendered from 12 views using randomly sampled cameras and lighting from a specific set of environmental lighting maps. To evaluate baselines and our method, we hold out 70 random meshes from Objaverse [10] as the test set.We additionally gather a collection of test meshes from 3D game assets to assess our method, showing its broad applicability. Further details can be found in the appendix.
@@
Figure 7: Qualitative Analysis. We visually compare our method with existing baselines. Although Fantasia3D [8] also attempts to generate relightable texture, unlike ours, their result still tends to have baked-in lighting. For non-relightable texture generation, we replace our LightControlNet with depth ControlNet and achieve superior results in terms of both image quality and runtime compared to the baselines. More details are in Table 1.
Objaverse FID $\downarrow$ KID $\downarrow$ Runtime $\downarrow$
subset (×$\times 10^{-3}$) (mins)
Latent-Paint [25] 73.65 7.26 10
Fantasia3D [8] 120.32 8.34 30
TEXTure [38] 71.64 5.43 6
Text2tex [7] 95.59 4.71 15
Ours (w/ depth) 60.49 3.96 2
Ours 62.67 2.69 4
Table 1: Quantitative Evaluation. We test our methods and baselines on 70 objects from Objaverse [10]. With depth ControlNet, our method yields superior results to all baselines in a rapid runtime of 2 minutes. Using LightControlNet (Ours) within our model improves the lighting disentanglement of our texture generation and maintains comparable image quality.
@@
Qualitative Analysis and User Study. As shown in Figure 6, our method can generate highly-detailed textures that can be rendered properly with the environment lighting across a wide variety of meshes. We also visually compare our method and the baselines in Figure 7. Our method produces textures with higher visual fidelity than the baselines for both the relightable and non-relightable variants. In particular, When compared with Fantasia3D [8], a recent work that also aims to generate material-based texture, our results not only have superior visual quality, but also disentangle the lighting more successfully. To further evaluate the quality of the relighted texture quantitatively, we conduct a user study asking participants to compare our results with Fantasia3D. Both results are rendered under identical lighting environments to ensure a fair evaluation, with a shaded mesh image in display signifying the lighting direction (please refer to the appendix for more details). 86.7% of the participants prefer our results, indicating that our approach consistently surpasses Fantasia3D in relighting quality.
Objaverse FID $\downarrow$ KID $\downarrow$ Runtime $\downarrow$
subset (×$\times 10^{-3}$) (mins)
@@
Table 2: Ablation Study. We perform an ablation analysis on our distilled encoder (1st row) and multi-view visual prompting (2nd row). Replacing the distilled encoder with the original VQ-VAE encoder doubled the time of performance without a noticeable quality improvement. When removing the multi-view visual prompting for initial generation, the system requires 2000 iterations (5x slowdown compared to our 400 iterations) to produce reasonable results, which leads to slightly worse texture quality.
Ablation Study. We perform an ablation analysis on our multi-view visual prompting (m.v.v.p.) and distilled encoder (dist. enc.) as seen in Table 2. When substituting our distilled encoder with the original VQ-VAE encoder, the performance is twice as slow, but the quality of results is not noticeably superior. On the other hand, without the multi-view visual prompting for the initial generation, the system requires 2000 iterations (5x slowdown compared to our 400 iterations) to produce reasonable results, while still leading to slightly worse texture quality.
@@ 6 Discussion
We proposed an automated texturing technique based on user-provided textual descriptions. Our method employs an illumination-aware 2D diffusion model (LightControlNet) and an improved optimization process based on the SDS loss. Our approach is substantially faster than previous methods while yielding high-fidelity textures with illumination disentangled from surface reflectance/albedo. We demonstrated the efficacy of our method through quantitative and qualitative evaluation on the Objaverse dataset.
@@
Acknowledgments. We thank Benjamin Akrish, Victor Zordan, Dmitry Trifonov, Derek Liu, Sheng-Yu Wang, Gaurav Parmer, Ruihan Gao, Nupur Kumari, and Sean Liu for their discussion and help. The project is partly supported by Roblox. JYZ is partly supported by the Packard Fellowship. KD is supported by the Microsoft Research PhD Fellowship.
Appendix A Implementation Details
Distilled Encoder. In Section 4.1 (main paper), we improve the efficiency of LightControlNet by distilling the image encoder in Stable Diffusion [39]. We profile the running time of our distilled and original encoder in the table below.
@@ Appendix A Implementation Details
Hyper-parameters. We provide the hyper-parameters used by our pipeline in Section 4.3 (main paper). $\lambda_{\text{MSE}}=1000,\lambda_{\text{VGG}}=1000,\lambda_{\text{Reg}}=10$. We use a batch of 4, a learning rate of 0.01 for optimization, and a CFG scale of 50 in Score Distillation Sampling loss. In Section 4.1 and Figure 4 (main paper), we set three pre-defined materials to generate a conditioning image. The specific material parameters are (1) non-metal, non-smooth: $k_{m}=0,k_{r}=1$; (2) half metal, half smooth: $k_{m}=0.5,k_{r}=0.5$; (3) pure metal, extremely smooth: $k_{m}=1,k_{r}=0$. The color $k_{c}$ is always set to $(1,1,1)$.
Base Models. We use stable diffusion v1.5 (SD1.5) as our base model for the experiments in Table 1, Table 2 and Figure 7 (main paper). Our pipeline is also compatible with other base models fine-tuned from SD1.5. For example, we have also used Dreamshaper, a community fine-tuned checkpoint of SD1.5, to generate a variety of captivating textures. We include some of the results in Figure 6 (main paper) and Figure 11. Specifically, the results of the jackets, goblins, fishmen, and wolves are obtained using Dreamshaper.
Environment Maps. In Section 4, we use randomly rotated environment maps to represent different lighting conditions. Specifically, we download 6 HDRI light maps from polyhaven. These HDRI maps are captured in a studio environment. We show these light maps in Figure 8. We use the middle left one as the fixed lighting $L^{*}$ in Section 4.2.
Figure 8: Environment Maps. We download 6 HDRI maps from polyhaven to represent different different lighting conditions. For random lighting samples, we select one map from them and apply a random rotation.
Evaluation and User Study. For quantitative evaluation in Table 1 and Table 2 (main paper), each resulting textured object is rendered from a fixed set of 16 surrounding viewpoints. FID and KID are computed against the ground-truth renderings of the same viewpoints using Objaverse’s [10] original texture. For the user study mentioned in Section 5 (main paper), we compare our method with Fantasia3D’s relightable texture [8]. Using both methods, we generate the texture for 10 test objects and place them in identical lighting conditions. We also render an untextured mesh under this light to illustrate the reference lighting direction, as depicted in Figure 12. We then ask 15 participants to identify the texture with greater fidelity in adhering to the reference lighting condition.
BRDF Model and Rendering Equation. As described in Section 4.3, our material model [48] consists of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ which is a perturbation of surface normal in tangent space, and the base color $k_{c}\in\mathbb{R}^{3}$. The general rendering equation is:
@@
Appendix B Additional Experiments
Backprojection Baseline. We mention an alternative baseline by directly backprojecting the sparse views in Section 4. We compare this baseline with our method in Figure 9.
@@ Appendix B Additional Experiments
Evaluation on Game-Asset Data. We collect 22 meshes from 3D game assets with 5 prompts each and evaluate our method and baselines on this new dataset. Table 4 shows our method outperforms the baselines on this new dataset.
3D Game Assets KID $\downarrow$
(×$\times 10^{-3}$)
Latent-Paint [25] 9.34
TEXTure [38] 5.64
Text2tex [7] 5.21
Ours 3.43
Table 4: Quantitative Evaluation on Game-Asset Data. We compare our method with the baselines on our newly collected dataset outside of Objaverse. The results show our method achieves better results than the baselines.
Ablation on material bases. In Section 4.1, we render a conditioning image using 3 pre-defined materials to encompass a broad range of feasible materials: (1) non-metal, not smooth (diffuse effect); (2) half-metal, half-smooth (mixed effect); (3) pure metal, smooth (specular effect). We present an ablation study of these material bases in Table 5:
Material Basis Ours:(1),(2),(3) w/o (1) w/o (2) w/o (3) w/o (1)(3)
FID $\downarrow$ 62.67 66.34 64.32 67.43 72.13
KID $\downarrow$ 2.69 3.11 3.42 4.12 4.53
Table 5: Ablation on material bases.
Additional Results. We show additional results in Figures 11 and 12. We also show an example of our generated material maps and GIF results with rotating lighting in Figure 10.
[…]
Figure 10: We show an example of our generated material maps in top row and GIF results with rotating lighting. Adobe Acrobat needs to be used to view the bottom results as GIFs.
Appendix C Societal Impact
Our method in text-based mesh texturing will enable many applications in 3D content creation. First, our method drastically reduces the time and expertise required for texture authoring, making 3D content generation accessible to a broader audience. Additionally, our faster inference time with improved result quality also reduces computational costs and energy consumption. However, our method might also be misused to generate fake content for misinformation. Nevertheless, humans can currently distinguish our synthesized objects from photo captures of real objects.
@@ Appendix C Societal Impact
Figure 12: Additional Qualitative Analysis. We present additional comparisons with the baselines. To evaluate the quality of relightable textures, we place our results and those from Fantasia3D under identical lighting conditions. An untextured mesh is also rendered under this condition to denote the reference lighting direction. Our results adhere to the lighting conditions and generally outperform all existing baselines in quality.

### v2

@@ 5 Experiments
Dataset. As illustrated in Figure 3, we employ Objaverse [11] to render paired data to train our LightControlNet. Objaverse consists of approximately 800k objects, of which we use the names and tags as their text descriptions. We filter out objects with low CLIP similarity [40] to their text descriptions and select around 40k as our training set. Each object is rendered from 12 views using randomly sampled cameras and lighting from a specific set of environmental lighting maps. To evaluate baselines and our method, we hold out 70 random meshes from Objaverse [11] as the test set. We additionally gather 22 mesh assets from 3D online games with 5 prompts each to assess our method, demonstrating its ability to generalize beyond Objaverse. Further details are in the appendix.
@@
Figure 7: Qualitative analysis. (a) We compare our method with Fantasia3D [9] that also attempts to generate Physically Based Rendering (PBR) texture. However, unlike ours, their results often exhibit baked-in lighting, leading to artifacts when put into varied lighting environments. (b) We also compare our method with other baselines that can only generate non-relightable (RGB) texture. For non-relightable texture generation, we can replace our LightControlNet with depth ControlNet and generate RGB textures with a shorter runtime. More details are in Table 1.
Table 1: Quantitative Evaluation. We test our methods and baselines on 70 test objects from Objaverse [11] and 22 objects curated from 3D game assets. With depth ControlNet, our method yields superior results to all baselines while being three times as fast as the fastest baseline. Using LightControlNet (Ours) within our model improves the lighting disentanglement while maintaining comparable image quality.
Objaverse test set Game Asset Runtime $\downarrow$
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
(×$\times 10^{-3}$) (×$\times 10^{-3}$) (mins)
Latent-Paint [29] 73.65 7.26 204.43 9.25 10
Fantasia3D [9] 120.32 8.34 164.32 9.34 30
TEXTure [43] 71.64 5.43 103.49 5.64 6
Text2tex [8] 95.59 4.71 119.98 5.21 15
Ours (w/ depth) 60.49 3.96 85.92 3.87 2
Ours 62.67 2.69 83.32 3.34 4
Table 2: User study. We conduct a user preference study to evaluate (1) result realism, (2) texture consistency with input text, and (3) plausibility under varied lighting. Participants consistently prefer our results over all baselines in these respects.
Preferred Percentage Objaverse test set
Realistic Consistent with text Relightable
Ours v.s. Latent-Paint [29] $92.6\%$ $74.5\%$ $84.3\%$
Ours v.s. Fantasia3D [9] $81.9\%$ $67.6\%$ $74.3\%$
Ours v.s. TEXTure [43] $70.8\%$ $57.3\%$ $87.1\%$
Ours v.s. Text2tex [8] $75.4\%$ $61.6\%$ $88.6\%$
@@
Qualitative analysis. As shown in Figure 6, our method can generate highly-detailed textures that can be rendered properly with the environment lighting across a wide variety of meshes. We also visually compare our method and the baselines in Figure 7. Our method produces textures with higher visual fidelity than the baselines for both the relightable and non-relightable variants. In particular, when compared with Fantasia3D [9], a recent work that also aims to generate material-based texture, our results not only have superior visual quality, but also disentangle the lighting more successfully.
User study. To further evaluate the texture quality quantitatively, we conduct a user study comparing our results with each of the baselines on the Objaverse test set in Table 2. We ask 30 participants to evaluate (1) the realism of the results, (2) the consistency of the generated texture with the input text, and (3) the plausibility of the results when placed under varying lighting conditions. Each result is presented in the form of 360-degree rotation to display full texture details. The reference lighting is provided alongside when participants evaluate (3). Across all three aspects, participants consistently prefer our method over baselines.
Table 3: Ablation study on algorithmic components. We analyze the role of our distilled encoder (1st row) and multi-view visual prompting (2nd row). Replacing the distilled encoder with the original VQ-VAE encoder doubles the running time without a noticeable improvement. When removing the multi-view visual prompting for initial generation, the system requires 2,000 iterations (5x compared to our 400 iterations) to produce reasonable results, which produces slightly worse texture quality.
Objaverse test set FID $\downarrow$ KID (×$\times 10^{-3}$) $\downarrow$ Runtime $\downarrow$ (mins)
@@
Table 4: Ablation study on material bases. We verify the impact of the material bases in rendering conditioning images. Omitting any one of these degrades quality.
Material Basis
non-metal, half-metal, pure metal, FID $\downarrow$ KID (×$\times 10^{-3}$) $\downarrow$
not smooth half-smooth smooth
62.67 2.69
66.34 3.11
64.32 3.42
67.43 4.12
72.13 4.53
Table 5: Ablation study on the number of canonical views. We analyze the role of our canonical view selection in Section 4.2. Relying on only the left and right views provides insufficient supervision. Interestingly, adding top and bottom views leads to worse overall quality. This is likely due to the limitation of pre-trained 2D diffusion models in synthesizing top and bottom views well for a variety of objects. Furthermore, given the fixed resolution of the multi-view image, stacking more views would result in a lower resolution for each view, leading to a worse initialization for Stage 2.
Num. of canonical views FID $\downarrow$ KID (×$\times 10^{-3}$) $\downarrow$
2 views (front, back) 67.43 3.47
4 views (Ours: front, back, left, right) 62.67 2.69
6 views (front, back, left, right, top, bottom) 70.14 3.72
Ablation study. We perform a thorough ablation analysis on different aspects of our method as seen in Table 3. When substituting our distilled encoder with the original VQ-VAE encoder, the performance is twice as slow, but the quality of results is not noticeably superior. On the other hand, without the multi-view visual prompting for the initial generation, the system requires 2000 iterations (a 5x slowdown compared to our 400 iterations) to produce reasonable results, while still leading to slightly worse texture quality.
In Section 4.1, we render a conditioning image using 3 pre-defined materials to encompass a broad range of feasible materials: (1) non-metal, not smooth (diffuse effect); (2) half-metal, half-smooth (mixed effect); (3) pure metal, smooth (specular effect). These material bases cover a large range of feasible materials. Table 4 shows omitting any one of these bases degrades quality.
As shown in Table 5, we also evaluate our selection of four canonical views in Section 4.2. Relying on only the left and right views provides insufficient supervision. Interestingly, incorporating top and bottom views degrades the performance. We hypothesize that the degradation is likely due to the limitation of 2D diffusion model backbones in generating top and bottom views reliably. Furthermore, stacking more views within a single image results in a decreased resolution for each view, given the fixed resolution of the multi-view image.
@@ 6 Discussion
We proposed an automated texturing technique based on user-provided prompts. Our method employs an illumination-aware 2D diffusion model (LightControlNet) and an improved optimization process based on the SDS loss. Our approach is substantially faster than previous methods while yielding high-fidelity textures with illumination disentangled from surface reflectance/albedo. We demonstrated the efficacy of our method through quantitative and qualitative evaluation on the Objaverse dataset and meshes curated from game assets.
@@
Appendix 0.A Implementation Details and Additional Results
Distilled Encoder. In Section 4.1 of the main paper, we improve the efficiency of LightControlNet by distilling the image encoder in Stable Diffusion [44]. We profile the running time of our distilled and original encoder in the following table.
@@ Appendix 0.A Implementation Details and Additional Results
Hyper-parameters. We provide the hyper-parameters used by our pipeline in Section 4.3 of the main paper. We use a batch of 4, a learning rate of 0.01 for optimization, and a CFG scale of 50 in Score Distillation Sampling loss. In Section 4.1 and Section 4.2 (main paper), we set 3 pre-defined materials to generate a conditioning image. The specific material parameters are (1) non-metal, non-smooth: $k_{m}=0,k_{r}=1$; (2) half metal, half smooth: $k_{m}=0.5,k_{r}=0.5$; (3) pure metal, extremely smooth: $k_{m}=1,k_{r}=0$. The color $k_{c}$ is always set to $(1,1,1)$. We train our LightControlNet for 20,000 iterations with a batch size of 16.
Base Models. We use stable diffusion v1.5 (SD1.5) as our base model for the experiments in Tables 1, 2, and 3 of the main paper. Our pipeline is also compatible with other base models fine-tuned from SD1.5. For example, we have also used Dreamshaper, a community fine-tuned checkpoint of SD1.5, to generate a variety of captivating textures. We include some of the results in Figure 6 (main paper) and Figure 10. Specifically, the results of the jackets, goblins, fishmen, and wolves are obtained using Dreamshaper.
Environmental Light Maps. In Section 4, we use randomly rotated environmental light maps to represent different lighting conditions. Specifically, we download 6 HDRI light maps from polyhaven. These HDRI maps are captured in a studio environment. We show these light maps in Figure 8. We also study the effect of different lighting by using various environment maps in training LightControlNet. We contrast the indoor set used in the main paper (FID: 62.67, KID: 2.69) with an outdoor set of 6 maps featuring brighter ambient light (FID: 63.23, KID: 2.66), and another set of 6 maps with less directional lighting (FID: 70.14, KID: 3.72). The results suggest directional lighting is important in accurately modeling lighting effects with LightControlNet, whereas ambient light intensity minimally impacts quality.
Figure 8: Environmental Light Maps. We download 6 HDRI maps from polyhaven to represent different different lighting conditions. For random lighting samples, we select one map from them and apply a random rotation.
BRDF Model and Rendering Equation. As described in Section 4.3 of the main paper, our material model [53] consists of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ which is a perturbation of surface normal in tangent space, and the base color $k_{c}\in\mathbb{R}^{3}$. We show an example of our generated material maps in Figure 12. The general rendering equation is:
@@
Rendering Engine. We use nvdiffrast [20], a differentiable renderer, as our rendering engine in both of our training and testing phases. Notably, our exported material maps are also directly compatible with widely used rendering applications, such as Blender. This compatibility is demonstrated in the main paper, where Figure 1 and the top half of Figure 7 are rendered using Blender, showcasing the practical application of our generated textures.
@@
Figure 11: Additional Qualitative Analysis. We present additional comparisons with the baselines. To evaluate the quality of relightable textures, our results and those from Fantasia3D are placed under identical lighting conditions. An untextured mesh is also rendered under this condition to denote the reference lighting direction. Our results adhere to the lighting conditions and generally outperform all existing baselines in quality.
Backprojection Baseline. We mention an alternative baseline by directly backprojecting the sparse views in Section 4 of the main paper. We compare this baseline with our method in Figure 9.
Additional Results. We show additional results in Figures 10 and 11. We also show an example of our generated material maps in Figure 12.
Figure 12: We show an example of our generated material maps.
Appendix 0.B Additional Related Works
In Section 2 of the main paper, we discuss prior works on 3D texture generation. Another excellent recent work, TexFusion [6], aligns with the research using depth-conditioned diffusion models for texture creation. They introduce a Sequential Interlaced Multiview Sampler, which performs denoising steps across multiple camera viewpoints and integrates the outcomes into a unified latent texture map. The execution time for TexFusion is comparable to that of TEXTure[43], with the comprehensive approach taking approximately 6 minutes, whereas a quicker variant takes about 2.2 minutes but yields lower-quality results. Similar to both TEXTure[43] and Text2tex[8], TexFusion’s textures are entangled with lighting effects. The non-relightable textures make it challenging to adapt to various lighting conditions. On the other hand, our generated texture can be relit properly in different lighting environments with a comparable runtime (4 mins). As their code is unavailable, we cannot preform a direct comparison. However, we will be happy to compare our method with this baseline once their code becomes available.
Appendix 0.C Societal Impact
Our method in text-based mesh texturing will enable many applications in 3D content creation. First, our method drastically reduces the time and expertise required for texture authoring, making 3D content generation accessible to a broader audience. Additionally, our faster inference time and improved result quality also reduce computational costs and energy consumption. However, our method might also be misused to generate fake content for misinformation. Nevertheless, we believe humans can currently distinguish our synthesized objects from photo captures of real objects.